# A Semantics-Aware Hierarchical Self-Supervised Approach to Classification of Remote Sensing Images

### v1

@@ Abstract
Deep learning has become increasingly important in remote sensing image classification due to its ability to extract semantic information from complex data. Classification tasks often include predefined label hierarchies that represent the semantic relationships among classes. However, these hierarchies are frequently overlooked, and most approaches focus only on fine-grained classification schemes. In this paper, we present a novel Semantics-Aware Hierarchical Consensus (SAHC) method for learning hierarchical features and relationships by integrating hierarchy-specific classification heads within a deep network architecture, each specialized in different degrees of class granularity. The proposed approach employs trainable hierarchy matrices, which guide the network through the learning of the hierarchical structure in a self-supervised manner. Furthermore, we introduce a hierarchical consensus mechanism to ensure consistent probability distributions across different hierarchical levels. This mechanism acts as a weighted ensemble being able to effectively leverage the inherent structure of the hierarchical classification task. The proposed SAHC method is evaluated on three benchmark datasets with different degrees of hierarchical complexity on different tasks, using distinct backbone architectures to effectively emphasize its adaptability. Experimental results show both the effectiveness of the proposed approach in guiding network learning and the robustness of the hierarchical consensus for remote sensing image classification tasks.
@@ I Introduction
Image classification and semantic segmentation are fundamental tasks in image processing. While Deep Learning (DL) methods have significantly advanced the field by learning powerful feature representations, standard approaches often treat classification as a flat, single-level task. However, many applications involve target categories having a natural hierarchical structure, exhibiting multiple levels of semantic granularity (e.g., object taxonomies, language, land-cover systems) [1, 2, 3]. Neglecting this inherent structure discards valuable semantic information and can lead to suboptimal or inconsistent predictions [4]. This is particularly relevant when dealing with label ambiguity or variations in image quality (such as resolution or artifacts) that may only permit reliable classification at coarser semantic levels.
A key challenge is integrating hierarchical information effectively into the learning process. Traditional methods frequently rely on post-processing steps or fixed rule-based systems to enforce hierarchical consistency, which limit their ability to learn complex, data-driven relationships between levels [5, 6, 7, 8]. Furthermore, constructing large-scale datasets with consistent multi-level annotations is often difficult [9, 10]. This requires methods capable of handling data labelled at various levels within the hierarchy. Although standardized hierarchical classification schemes exist in several domains, e.g., COoRdination of INformation on the Environment (CORINE) and Land Cover Classification System (LCCS) in Remote Sensing (RS), most current DL models do not explicitly leverage this structure during training.
Most hierarchical approaches focus on extracting features at multiple levels of granularity without explicitly modeling the hierarchical relationships intrinsic to the task [11, 12]. This poses a limitation, as the hierarchy arises from the problem domain itself rather than from specific architectural design choices. Accordingly, an effective hierarchical learning principle should be generalizable and scalable across different DL architectures. However, existing methods often produce non-modular architectures with rigid structures, thereby limiting the flexibility in the DL backbone selection.
In this paper, we propose a novel image processing methodology to modularly improve scene classification and semantic segmentation through semantic relationships across categories. Our approach leverages the hierarchical structure of standardized classification schemes by introducing a classification head for each hierarchical level. We further introduce trainable hierarchy matrices to model the relationships between these levels. To enforce consistency across the hierarchy, we define a multi-level consensus prediction that promotes agreement among the predictions at different levels. Additionally, we introduce a self-supervised consensus loss that implicitly learns the hierarchical relationships within the data. This approach allows the discovery of intricate sematic relationships that can deviate from the simple user-defined aggregation rules, leading to improved classification accuracy and semantic consistency. We demonstrate the effectiveness of our approach on different RS image classification tasks, including semantic segmentation of Satellite Image Time Series (SITS) of multispectral acquisitions and Very-High Resolution (VHR) RS scene classification tasks.
@@
    We define a hierarchical image processing approach to semantic segmentation and scene classification that integrates multi-level classification modules with DL backbones, along with trainable projection matrices to model hierarchical semantic relationships.
@@
    We propose a novel self-supervised consensus loss function that enforces consistency across different hierarchical levels, enhancing the semantic coherence of the classification results.
@@
    The proposed method is general and can be applied to any DL architecture and hierarchical classification scheme, offering flexibility in defining architecture details.
@@
    We demonstrate the effectiveness of the proposed approach on three RS datasets that include images with varying resolutions and different complexities in their hierarchical relationships.
@@ II Related Works
Image classification is a fundamental task in computer vision and image processing, with DL models achieving state-of-the-art performance across various domains. Standard approaches typically treat classification as a “flat” problem in a traditional single-granularity classification, predicting probabilities over a set of mutually exclusive classes without explicitly considering potential semantic relationships or hierarchical structures among them. However, many real-world datasets possess inherent taxonomies where classes can be organized into hierarchies (e.g., species classification, object recognition with super-/sub-categories). Ignoring this structure can lead to suboptimal performance, as the model is not able to leverage prior knowledge about class relationships. Consequently, standard loss functions inherently treat all misclassifications equally. Without an explicit hierarchical representation, they fail to differentiate between errors involving closely related classes and those involving semantically distant ones.
@@
Leveraging hierarchical label structures during model training and inference has emerged as a promising direction to improve accuracy and semantic understanding. Early work demonstrated that standard Convolutional Neural Networks implicitly capture some hierarchical information, with shallower layers grouping high-level categories and deeper layers specializing in finer distinctions [13]. Building on this, researchers have explored explicit methods to incorporate hierarchy. Typically, this is achieved by mapping the label hierarchy to network architectures or by introducing loss functions that handle the hierarchical correlation between different granularity tasks [14, 15, 16]. Bertinetto et al. [17] proposed a level-wise hierarchical cross-entropy loss based on conditional probabilities within the hierarchy and used mapping functions to embed class relationships. Similarly, Garg et al. [18] introduced the Hierarchy Aware Features (HAF), incorporating multiple losses, including minimizing the Jensen-Shannon Divergence (JSD) between aggregated predictions at different hierarchy levels and enforcing geometric consistency in the feature space, aiming to reduce the severity of misclassifications. In [19], the authors proposed hierarchy and exclusion graphs to model semantic relations between two labels assigned to the same objects. They considered exclusions, overlaps, and subsumptions, and provided a detailed theoretical analysis of the structure of the hierarchical problem. Building on this work, a combinatorial loss is proposed in [20], which addresses the scenario where samples might be labelled at different levels of granularity. The loss is designed to maximize the marginal probability of the observed ground truth label by aggregating information from related labels in the hierarchy. The authors introduced a Hierarchical Residual Network (HRN) architecture, where features from parent levels act as residual connections to inform child-level features, explicitly facilitating hierarchical feature interaction.
@@
While these methods demonstrate the benefits of hierarchy-aware learning, they often rely on the assumption that a predefined and accurate hierarchy is available, where no hierarchy violation is possible. This assumption may not hold in all scenarios, as real-world taxonomies can be noisy, incomplete, or subjective. Furthermore, many existing techniques require explicit guidance or architectural modifications tailored to a specific fixed hierarchy [21], potentially limiting their flexibility and ability to adapt if the underlying hierarchical understanding evolves or is uncertain. For example, some approaches decompose the problem into sequential decisions or binary classifications guided by the known structure [22, 23, 21], rather than allowing the model to learn or refine hierarchical relationships implicitly from the data. The challenge remains to develop methods that can effectively exploit hierarchical semantics, even when the hierarchy is imperfect, labels are provided at varying granularity levels, or needs to be partially inferred in a robust and data-driven manner.
@@ III Proposed Methodology
In this work, we address the above-mentioned challenges by proposing a novel deep learning framework for hierarchy-aware image classification. Our approach models the hierarchical semantic relationships within image data by integrating hierarchical loss components with a robust consensus mechanism. This allows the model to leverage potential hierarchical structures effectively while mitigating the impact of potential inaccuracies or inconsistencies in the predefined taxonomy. Specifically, our method aims to learn and exploit hierarchical relationships in a more implicit, self-supervised fashion within a hierarchy-aware architecture, reducing the reliance on rigid, externally defined structure.
@@ III-A Hierarchical Problem Formulation
Let us denote column vectors by lowercase bold letters (e.g., $\mathbf{v}$) and matrices by uppercase bold letters (e.g., $\mathbf{B}$). The $i$th component of a vector $\mathbf{v}$ is denoted by $v_{i}$, while the $i$th row and $j$th column of a matrix $\mathbf{B}$ are denoted by $\mathbf{B}_{i,\cdot}$ and $\mathbf{B}_{\cdot,j}$, respectively. Then, let us consider a user-defined hierarchy label tree with $H$ levels, where the leaves represent the finest level of detail. We denote each hierarchical level as $h\in[1,2,\dots,H]$, ranging from the coarsest level $h=1$ to the finest one $h=H$. The root node $h=0$ is not considered, as it represents the trivial case in which all classes are mapped into a single class. Let $\mathbf{x}\in{\mathbb{R}^{b}}$ be the input vector, where $b$ is the number of input features. Its associated labels are $y^{h}\in\Omega^{h}$, where $\Omega^{h}=\{\omega^{h}_{1},\dots,\omega^{h}_{|\Omega^{h}|}\}$ is the set of labels at hierarchical level $h$, with $|\Omega^{h}|$ number of labels. For every $h_{1}<h_{2}$, we have that each fine-grained class $\omega_{j}^{h_{2}}$ is mapped to a coarse-grained class $\omega_{i}^{h_{1}}$, and $|\Omega^{h_{1}}|<|\Omega^{h_{2}}|$. We define this mapping relationship by $\omega_{j}^{h_{2}}\implies\omega_{i}^{h_{1}}$. Let $\mathcal{D}=\{(\mathbf{x}_{n},Y_{n})|n=1,\dots,N\}$ be the training set, where for each input feature vector $\mathbf{x}_{n}$, we have an associated set $Y_{n}=\{y_{n}^{1},\dots,y_{n}^{h},\dots,y_{n}^{H}\}$, which denotes the ground truth labels at all $H$ hierarchical levels, for a total number of $N$ training samples. The labels in the dataset satisfy a nested label structure, where each class at a finer level is contained within a class at a coarser level.
@@ III-B Fully-Supervised Multi-Level Learning
Figure 1: Overview of the proposed hierarchical structure within the backbone (left) and associated hierarchical label tree (right). Different classification modules are trained to classify different hierarchical levels from the features extracted by the backbone.
@@ III-C Hierarchy Matrices
The CCE loss functions at each hierarchical level lead to independent classifications, which are not constrained to follow the hierarchy definition formulated in the problem statement. To impose the hierarchical constraints, additional consensus predictions and self-supervised penalty terms are introduced. These penalty terms are defined through hierarchy projection matrices designed to map the correct relationships between the different layers.
Let us consider for simplicity the case of $H=2$ levels, with $h_{1}$ representing the coarse-grained labels and $h_{2}$ the fine-grained labels, i.e., $h_{1}<h_{2}$. We refer to $\Omega^{h_{1}}$ as the set of the macroclasses (i.e., the coarse-grained classes) and to $\Omega^{h_{2}}$ as the set of the microclasses (i.e., the fine-grained classes). We can define the hierarchy indicator matrix $\mathbf{I}^{h_{2},h_{1}}\in\{0,1\}^{|\Omega^{h_{2}}|\times|\Omega^{h_{1}}|}$, which represents the hierarchical relationship between level $h_{2}$ and $h_{1}$:
@@
This approach can be easily extended to a multi-level scenario, considering any given number $H$ of hierarchical levels. In this case, the total number of hierarchy matrices that can be determined is $((H-1)\cdot H)/2$, as also mappings of distant relationship levels can be defined (e.g., from the root of the label tree to the leaves, skipping intermediate levels) if needed. An example is illustrated in Fig. 2.
@@
The hierarchy indicator matrices are initialized with ones and zeros representing the user-defined mapping. Consequently, $I^{h_{2},h_{1}}_{i,j}$ assumes a value of one when the microclass $\omega_{i}^{h_{2}}$ is contained in the macroclass $\omega_{j}^{h_{1}}$ through the relationship $\omega_{i}^{h_{2}}\implies\omega_{j}^{h_{1}}$}, thus defining the matrix elements as follows:
$I^{h_{2},h_{1}}_{i,j}=\begin{cases}1&\text{if }{\omega_{i}^{h_{2}}\implies\omega_{j}^{h_{1}}}\\ 0&\text{otherwise}\end{cases}.$ (6)
It is important to note that, in the case of hard-hierarchies, each microclass corresponds to exactly one macroclass. However, this may not hold true as there might exist semantic overlaps in the hierarchy structure. Additionally, a distinction between user-defined aggregations and data-driven aggregations has to be made. While the first is defined a-priori for the problem at hand by the user, the latter is not known and must be inferred directly from the data, reflecting the true underlying structure of the problem considered. To take this into account, we can soften the hierarchical relationships by allowing the matrices to take any value between 0 and $1$. In this way, one microclass may be associated with multiple macroclasses with different confidences reflecting the semantic overlaps. To this purpose, we can define hierarchy joint distribution matrices $\mathbf{J}^{h_{2},h_{1}}\in{[0,1]}^{|\Omega^{h_{2}}|\times|\Omega^{h_{1}}|}$. Each term of the matrix $J^{h_{2},h_{1}}_{i,j}=P(\omega^{h_{1}}_{j},\omega^{h_{2}}_{i})$ models the probability of observing a relationship between microclass $\omega^{h_{2}}$ and macroclass $\omega^{h_{1}}$. This constraint relaxation allows us to better combine the user-defined aggregation problem with a data-driven centered architecture, by learning these matrices directly from the data.
III-D Self-Supervised Hierarchical Consensus Learning

### v2

# A Hierarchical Self-Consistent Regularization Approach to Satellite Image Time Series Classification

@@ Abstract
Deep learning has become increasingly important in remote sensing image classification due to its ability to extract semantic information from complex data. Classification tasks often include predefined label hierarchies that represent the semantic relationships among classes. However, these hierarchies are frequently overlooked, and most approaches focus only on fine-grained classification schemes. In this paper, we present a novel Semantics-Aware Hierarchical Consensus (SAHC) approach to learn hierarchical features and relationships by integrating hierarchy-specific classification heads within a deep network architecture, each specialized in different degrees of class granularity. The proposed approach employs trainable hierarchy matrices, which guide the network through the learning of the hierarchical structure in a self-consistent manner. Furthermore, we introduce a hierarchical consensus mechanism to ensure aligned probability distributions across different hierarchical levels. This mechanism acts as a weighted ensemble being able to effectively leverage the inherent structure of the hierarchical classification task. The proposed SAHC method is evaluated on two benchmark datasets with different degrees of hierarchical complexity on different tasks, considering varying spectral and spatial resolutions. Experimental results show both the effectiveness of the proposed approach in guiding network learning and the robustness of the hierarchical consensus for remote sensing image classification tasks. The codes will be released at https://github.com/rslab-unitrento/sahc.
@@ I Introduction
Image classification and semantic segmentation are central tasks in Remote Sensing (RS) image processing. In recent years, the proliferation of RS data and open distribution policies have significantly increased the availability of Land-Use and Land-Cover (LULC) products, along with their corresponding classification schemes [27]. These products play a critical role in various applications and are widely used by government agencies, municipalities, and ministries. However, global and regional LULC maps often differ in terms of resolution and are affected by semantic inconsistencies in class definitions and production algorithms [28]. While Deep Learning (DL) methods have revolutionized LULC mapping by capturing complex spatial and spectral dependencies, standard approaches typically treat classification as a flat, single-level task. In practice, Land-Cover (LC) classes are naturally organized into hierarchical label structures that reflect intrinsic levels of semantic granularity. This multi-level hierarchy is a property of the physical landscape. For instance, within a single dataset, a specific target might be unambiguously identified at a coarse-grained level (e.g., forest), while its finer-grained classification (e.g., deciduous versus coniferous) remains uncertain due to inherent label ambiguity, complex scene compositions, or the physical limits of the sensor. Neglecting this inherent hierarchical structure discards valuable semantic information and can lead to suboptimal or inconsistent predictions, particularly when dealing with label ambiguity or complex RS scenes. An example of this hierarchical problem definition can be seen in Fig. 1.
Figure 1: Hierarchical classification task example. Level 4 represents the coarsest class definition, where each class is progressively broken down into its fine-grained representations.
In the literature, most of the works related to RS image classification focus on single-level approaches, discarding the hierarchical information contained in the data. Hierarchical classification schemes, such as the CORINE Land Cover (CLC) maps, have mostly been adopted in object-based image classification methods based on Very-High Resolution (VHR) satellite images [3], often considering a limited number of classes to simplify the classification task. To validate the hierarchical structure, few works analyzed the inter-level relationships, considering decision trees in an object-based classification scenarios [25], or validation approaches based on intercomparison of different maps [23]. Instead, the wide majority of works in the RS domain focus on single-level approaches, discarding the hierarchical information contained in the data. While these methods show promising results at individual hierarchical levels, they fail to incorporate the hierarchical structure of the data into the decision-making process, overlooking important information included in the source data.
Only recently, hierarchical and multi-granularity classification methods have emerged to address these challenges in RS. These methods leverage the hierarchical structure to encode semantic relationships among labels, enabling knowledge transfer across different levels of granularity. Recent studies have demonstrated the benefits of adaptive label granularity selection to balance classification accuracy and specificity [7], as well as the integration of parent-child relationships and sibling mutual exclusivity to improve hierarchical representation learning [9]. Despite these advancements, integrating hierarchical information effectively into the learning process remains an open challenge. Many existing methods rely on complex, task-specific architectures or rigid rule-based post-processing steps, which limit their flexibility across different DL backbones and fail to explicitly enforce probability consistency across different hierarchical levels in a generalized manner.
To address these limitations, we propose a novel and modular Semantics-Aware Hierarchical Consensus method for learning hierarchical features and relationships. Our approach integrates hierarchy-specific classification heads within standard DL architectures, each specialized in different degrees of class granularity. By introducing trainable projection matrices and explicitly modeling the semantic relationships between these levels, the proposed method ensures consistent predictions across the hierarchy without relying on rigid aggregation rules.
@@
    We define a hierarchical approach to semantic segmentation and scene classification that integrates multi-level classification modules with DL backbones, along with trainable projection matrices to model hierarchical semantic relationships.
@@
    We propose a novel self-consistent consensus loss function that enforces consistency across different hierarchical levels, improving the semantic coherence of the classification results.
@@
    The proposed method is general and can be applied to any DL backbone and hierarchical classification scheme, offering flexibility in defining architecture details.
@@
    We demonstrate the effectiveness of the proposed approach on two RS datasets that include images with varying resolutions and different complexities in their hierarchical relationships.
@@ II Related Works
Image classification and semantic segmentation are widely studied tasks in RS image processing, with DL models achieving state-of-the-art performance across various domains. Standard approaches typically treat classification as a “flat”, single-granularity problem, predicting probabilities over a set of mutually exclusive classes without explicitly considering potential semantic relationships or hierarchical structures among them. However, many real-world RS datasets possess inherent taxonomies where classes can be naturally organized into hierarchies (e.g., standardized land-cover systems, target recognition taxonomies with super- and sub-categories). Ignoring this structure can lead to suboptimal performance, as standard loss functions treat all misclassifications equally. Without an explicit hierarchical representation, model fail to differentiate between errors involving closely related classes and those involving semantically distant ones.
@@
Leveraging hierarchical label structures during model training and inference has emerged as a promising direction to improve accuracy and semantic understanding. Early work demonstrated that standard Convolutional Neural Networks implicitly capture some hierarchical information, with shallower layers grouping high-level categories and deeper layers specializing in finer distinctions [2]. Building on this, researchers have explored explicit methods to incorporate hierarchy. Typically, this is achieved by mapping the label hierarchy to specific network architectures or by introducing loss functions that handle the correlation between different granularity tasks [5, 14, 17]. For instance, several methods propose level-wise hierarchical cross-entropy losses based on conditional probabilities, or introduce custom features to minimize the divergence between aggregated predictions at different hierarchical levels, aiming to reduce the severity of misclassification. Bertinetto et al. [1] proposed a level-wise hierarchical cross-entropy loss based on conditional probabilities within the hierarchy and used mapping functions to embed class relationships. Similarly, Garg et al. [15] introduced the Hierarchy Aware Features (HAF), incorporating multiple losses, including minimizing the Jensen-Shannon Divergence (JSD) between aggregated predictions at different hierarchy levels and enforcing geometric consistency in the feature space, aiming to reduce the severity of misclassifications. In [12], the authors proposed hierarchy and exclusion graphs to model semantic relations between two labels assigned to the same objects. They considered exclusions, overlaps, and subsumptions, and provided a detailed theoretical analysis of the structure of the hierarchical problem. Building on this work, a combinatorial loss is proposed by Chen et al. [8], which addresses the scenario where samples might be labelled at different levels of granularity. The loss is designed to maximize the marginal probability of the observed ground truth label by aggregating information from related labels in the hierarchy. The authors introduced a Hierarchical Residual Network (HRN) architecture, where features from parent levels act as residual connections to inform child-level features, explicitly facilitating hierarchical feature interaction. Following this work, Zhou et al. [31] proposed a multi-granularity archaeological dating method that depends on an explicit knowledge-guided relation graph, employing a focal-type probability classification loss to mine the relationship between specific object attributes and their broader eras. Their network leverages a bidirectional multi-granularity module with gradient truncated addition to explicitly fuse coarse- and fine-grained features. Similarly, Cai et al. [4] introduced HierSwin, a hierarchical vision transformer, which explicitly models the semantic correlations embedded within a predefined three-level hierarchical tree to gradually improve accuracy from fine- to coarse-grained classifications.
More recently, some works addressed the Multigranularity Classification (MGC) in RS to handle targets that are discernible only at specific variable semantic levels due to differences in spatial resolution, sensor modalities, and data quality. Chen et al. [7] introduced an accuracy-specificity curve to adaptively select task-dependent prediction thresholds. Their approach also employs a penalty term based on the global semantic distance across the label tree to adjust class logits. Addressing a similar challenge in ship classification, Chen et al. [9] proposed a hierarchical contrastive learning framework. By explicitly incorporating parent-child dependencies and mutual exclusivity among sibling categories, their method aligns class-specific representations with semantic relationships and enforces probability coherence across levels through a consistency loss.
@@
Despite these significant advancements, many existing multigranularity and hierarchical learning methods still rely on the assumption that a predefined and accurate hierarchy is available, where no hierarchy violation is possible. This requires rigid post-processing rules, fixed decision thresholds, or highly specific non-modular architectures. Furthermore, many existing techniques require explicit guidance or architectural modifications tailored to a specific fixed hierarchy [18], potentially limiting their flexibility and ability to adapt if the underlying hierarchical understanding evolves or is uncertain. For example, some approaches decompose the problem into sequential decisions or binary classifications guided by the known structure [11, 29], rather than allowing the model to learn or refine hierarchical relationships implicitly from the data. The challenge remains to develop methods that can effectively exploit hierarchical semantics, even when the hierarchy is imperfect, labels are provided at varying granularity levels, or needs to be partially inferred in a robust and data-driven manner.
@@ III Proposed Methodology
In this work, we address the above-mentioned challenges by proposing a novel deep learning framework for hierarchy-aware image classification. Our approach models the hierarchical semantic relationships within image data by integrating hierarchical loss components with a robust consensus mechanism. This allows the model to leverage potential hierarchical structures effectively while mitigating the impact of potential inaccuracies or inconsistencies in the predefined taxonomy. Specifically, our method aims to learn and exploit hierarchical relationships in a more implicit, self-consistent fashion within a hierarchy-aware architecture, reducing the reliance on rigid, externally defined structure.
@@ III-A Hierarchical Problem Formulation
Let us denote column vectors by lowercase bold letters (e.g., $\mathbf{v}$) and matrices by uppercase bold letters (e.g., $\mathbf{B}$). The $i$th component of a vector $\mathbf{v}$ is denoted by $v_{i}$, while the $i$th row and $j$th column of a matrix $\mathbf{B}$ are denoted by $\mathbf{B}_{i,\cdot}$ and $\mathbf{B}_{\cdot,j}$, respectively. Then, let us consider a user-defined hierarchy label tree with $H$ levels, where the leaves represent the finest level of detail. We denote each hierarchical level as $h\in[1,2,\dots,H]$, ranging from the coarsest level $h=1$ to the finest one $h=H$. The root node $h=0$ is not considered, as it represents the trivial case in which all classes are mapped into a single class. Let $\mathbf{x}\in{\mathbb{R}^{b}}$ be the input vector, where $b$ is the number of input features. Its associated labels are $y^{h}\in\Omega^{h}$, where $\Omega^{h}=\{\omega^{h}_{1},\dots,\omega^{h}_{|\Omega^{h}|}\}$ is the set of labels at hierarchical level $h$, with $|\Omega^{h}|$ number of labels. For every $h_{1}<h_{2}$, we have that each fine-grained class $\omega_{j}^{h_{2}}$ is mapped to a coarse-grained class $\omega_{i}^{h_{1}}$, and $|\Omega^{h_{1}}|\leq|\Omega^{h_{2}}|$. We define this mapping relationship by $\omega_{j}^{h_{2}}\implies\omega_{i}^{h_{1}}$. Let $\mathcal{D}=\{(\mathbf{x}_{n},Y_{n})|n=1,\dots,N\}$ be the training set, where for each input feature vector $\mathbf{x}_{n}$, we have an associated set $Y_{n}=\{y_{n}^{1},\dots,y_{n}^{h},\dots,y_{n}^{H}\}$, which denotes the ground truth labels at all $H$ hierarchical levels, for a total number of $N$ training samples. The labels in the dataset satisfy a nested label structure, where each class at a finer level is contained within a class at a coarser level.
@@ III-B Fully-Supervised Multi-Level Learning
Figure 2: Illustration of the proposed hierarchical architecture with the hierarchical consensus calculated on $h=\tilde{h}$. Heads tailored to different granularity levels are reprojected on the target level and averaged to extract the consensus probability $\hat{p}^{\tilde{h}}_{HC}$. Other than for prediction, this consensus probability is used to calculate the hierarchical self-consistency loss $\mathcal{L}^{h}_{HC}$.
@@ III-C Hierarchy Matrices
The CCE loss functions at each hierarchical level lead to independent classifications, which are not constrained to follow the hierarchy definition formulated in the problem statement. To impose the hierarchical constraints, additional consensus predictions and self-consistent penalty terms are introduced. These penalty terms are defined through hierarchy projection matrices designed to map the correct relationships between the different layers.
Let us consider for simplicity the case of $H=2$ levels, with $h_{1}$ representing the coarse-grained labels and $h_{2}$ the fine-grained labels, i.e., $h_{1}<h_{2}$. We refer to $\Omega^{h_{1}}$ as the set of the macroclasses (i.e., the coarse-grained classes) and to $\Omega^{h_{2}}$ as the set of the microclasses (i.e., the fine-grained classes). We can define the hierarchy indicator matrix $\mathbf{I}^{h_{2},h_{1}}\in\{0,1\}^{|\Omega^{h_{2}}|\times|\Omega^{h_{1}}|}$, which represents the user-defined hierarchical relationship between level $h_{2}$ and $h_{1}$:
@@
This approach can be easily extended to a multi-level scenario, considering any given number $H$ of hierarchical levels. In this case, the total number of hierarchy matrices that can be determined is $((H-1)\cdot H)/2$, as also mappings of distant relationship levels can be defined (e.g., from the root of the label tree to the leaves, skipping intermediate levels) if needed. A generalized example is illustrated in Fig. 3.
@@
In a strict hard-hierarchy, $I^{h_{2},h_{1}}_{i,j}$ assumes a value of one when the microclass $\omega_{i}^{h_{2}}$ is contained in the macroclass $\omega_{j}^{h_{1}}$ through the relationship $\omega_{i}^{h_{2}}\implies\omega_{j}^{h_{1}}$}, and zero otherwise. However, this rigid formulation often fails to account for semantic overlaps in real-world hierarchies. Therefore, we distinguish between user-defined aggregations, established a priori for a given task, and data-driven aggregations, which are inferred directly to capture the data’s true underlying organization. To combine the user-defined prior with a data-driven optimization, we parameterize the hierarchical relationships in the unconstrained logit space. Instead of stricly fixing the matrices to binary values or directly bounding them in $[0,1]$, we define learnable parameter matrices $\mathbf{W}^{h_{2},h_{1}}$. These matrices are initialized based on the user-defined indicator matrix $\mathbf{I}^{h_{2},h_{1}}$ to provide a stronger prior:
$W^{h_{2},h_{1}}_{i,j}=\begin{cases}0.5+\epsilon&\text{if }{I^{h{2},h_{1}}_{i,j}=1}\\ -5.0+\epsilon&\text{if }{I^{h{2},h_{1}}_{i,j}=0}\end{cases},$ (6)
where $\epsilon\sim\mathcal{N}(0,0.01^{2})$ is a small random noise added to break network symmetry during initial phases of optimization. The two selected values act as positive and negative biases towards valid connections and unmapped relationships. During the forward pass, these unconstrained parameters are dynamically mapped into valid hierarchy log joint distribution matrices $\mathbf{L}^{h_{2},h_{1}}=\log{\mathbf{J}^{h_{2},h_{1}}}\in{\mathbb{R}}^{|\Omega^{h_{2}}|\times|\Omega^{h_{1}}|}$ via a log-softmax operation on the flattened parameter matrices $\mathbf{W}^{h_{2},h_{1}}$. Each term of the matrix $J^{h_{2},h_{1}}_{i,j}=P(\omega^{h_{1}}_{j},\omega^{h_{2}}_{i})$ models the probability of observing a relationship between microclass $\omega^{h_{2}}$ and macroclass $\omega^{h_{1}}$. This constraint relaxation allows us to better combine the user-defined aggregation with a data-driven centered architecture, by learning these matrices directly from the data while still reliably considering the initial user-defined mappings.
III-D Self-Consistent Hierarchical Consensus Learning